### v1

# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

Abstract
We introduce Sundial, a family of native, flexible, and scalable time series foundation models. To predict the next-patch’s distribution, we propose a TimeFlow Loss based on flow-matching, which facilitates native pre-training of Transformers on time series without discrete tokenization. Conditioned on arbitrary-length time series, our model is pre-trained without specifying any prior distribution and can generate multiple probable predictions, achieving flexibility in representation learning beyond using parametric densities. Towards time series foundation models, we leverage minimal but crucial adaptations of Transformers and curate TimeBench with 1 trillion time points, comprising mostly real-world datasets and synthetic data. By mitigating mode collapse through TimeFlow Loss, we pre-train a family of Sundial models on TimeBench, which exhibit unprecedented model capacity and generalization performance on zero-shot forecasting. In addition to presenting good scaling behavior, Sundial achieves new state-of-the-art on both point forecasting and probabilistic forecasting benchmarks. We believe that Sundial’s pioneering generative paradigm will facilitate a wide variety of forecasting scenarios.
time series, foundation models, pre-training, Transformers
1 Introduction
Time series forecasting has fascinated people for thousands of years. Although people have been able to determine the time using instruments like sundials in 3000 BC, time series forecasting is intrinsically non-deterministic (Box et al., 2015). Therefore, generating the range of probable predictions is crucial for decision-making. The growing demand has facilitated numerous statistical approaches over the past decades (Hyndman, 2018; Box, 2013), which provide high-profile theories and probabilistic models to make reliable schedules. Recent advancements bring the boom of deftly designed models that automatically learn intricate dynamics and correlations from raw data (Oreshkin et al., 2019; Nie et al., 2022; Zhang & Yan, 2023; Liu et al., 2023). Despite the impressive performance, deep models necessitate task-specific training on sufficient in-distribution data. Motivated by advances in large models (Bommasani et al., 2021), pre-trained time series foundation models have shown promising capabilities in out-of-distribution tasks (Das et al., 2023b; Liu et al., 2024b; Woo et al., 2024; Ansari et al., 2024).
Figure 1: A native time series model learns the representation on the continuous token. A flexible foundation model is pre-trained without specifying the prior distribution. Sundial is presented as the first family of native and flexible time series foundation models.
Current research of time series foundation models has converged on building unified, scalable, and out-of-the-box forecasters, exhibiting zero-shot performance close to or sometimes surpassing supervised methods (Aksu et al., 2024). Notably, Transformers (Radford et al., 2018) are currently the de facto architecture of these models. While pre-trained Transformers with inherent generative capability have facilitated great success in language, image, and video generation (Ramesh et al., 2021; OpenAI, 2023; Liu et al., 2024c), most time series foundation models are not “generative” or, more specifically, probabilistic forecasters, thereby limiting reliability in decision-making. Although parametric densities specified with prior distributions (Wen et al., 2017; Woo et al., 2024) can be incorporated to cope with the uncertainty in predictions, they can constrain the capacity of distributions represented by a foundation model, especially on time series corpora characterized by high heterogeneity. To learn arbitrarily-intricate distributions without mode collapse, language modeling (Bengio et al., 2000) that learns the categorical distribution via cross-entropy loss inspires subsequent works (Gruver et al., 2023; Ansari et al., 2024), which treat time series as a foreign language using discrete tokenization. Still, apparent distinctions between continuous-valued time series and language tokens can lead to out-of-vocabulary issues and coarse-grained prediction intervals.
As shown in Figure 1, we leverage the native patching for time series tokenization and generative modeling to learn flexible predictive distributions for the first time. As foundation models aim to learn complicated distributions from extensive datasets and facilitate transferability across agnostic downstream tasks, we do not adopt any specific probabilistic priors, such as unimodal Gaussian or multimodal mixtures. To tame Transformers as native and scalable time series foundation models, we adopt generative modeling, a well-established paradigm that offers as much flexibility as language modeling and is better suited for continuous-value time series. Instead of adopting prevailing denoising diffusion models (Ho et al., 2020), we adopt a simple yet effective flow-matching framework (Lipman et al., 2022), which provides better efficiency and quality for generation (Tong et al., 2023). We propose TimeFlow Loss, formulated as a parameterized training objective, for autoregressive models to learn each token’s predictive distribution. This optimization objective accurately operates on original values and facilitates patch-level generation for quick inference, which is highly compatible with continuous-valued modalities.
Besides the TimeFlow Loss, we enhance Transformers with minimal but critical adaptations. We design patch embedding that is compatible with non-divisible context length. We adopt RoPE (Su et al., 2024) to enhance temporal causality. We leverage Pre-LN (Xiong et al., 2020), FlashAttention (Dao et al., 2022), and KV Cache (Pope et al., 2023), which are crucial but generally neglected in the development of time series foundation models. With the TimeFlow Loss facilitating the right paradigm for training scalable foundation models, these adaptations further optimize deployment.
To explore the scaling law of time series foundation models, we collect and curate TimeBench with an unprecedented volume of 1 trillion time points. We present Sundial as a family of highly capable foundation models, which achieve state-of-the-art on three large-scale and best-recognized benchmarks, including Time-Series-Library (TSLib) (Wu et al., 2022), GIFT-Eval (Aksu et al., 2024) and FEV (Ansari et al., 2024). Our contributions lie in these aspects:
  • We propose TimeFlow Loss to predict next-patch’s distribution, allowing Transformers to be trained without discrete tokenization and make probable predictions.
  • We present Sundial, a family of scalable and efficient time series foundation models pre-trained on 1 trillion time points, utilizing our enhanced Transformer.
  • Experimentally, Sundial achieves state-of-the-art zero-shot performance on point forecasting benchmarks and probabilistic forecasting leaderboards, including GIFT-Eval and FEV, positioning generative time series foundation models as a capable tool for decision-making.
2 Related Work
2.1 Time Series Forecasting
Forecasting is essential for decision-making, which has facilitated the development of statistical and deep-learning models. Advancements in deep models for time series include theory-inspired components (Wu et al., 2021; Zeng et al., 2023; Wu et al., 2022), architecture-oriented adaptations (Bai et al., 2018; Salinas et al., 2020; Lim et al., 2021), and time series processing (Kim et al., 2021; Nie et al., 2022). While deep models learning the dataset-level distribution enjoy large model capacity, statistical methods fitting separate dynamics of each time series are still prevailing choices due to their flexibility and performance on small data (Ke et al., 2017; Hyndman, 2018).
One of the efforts towards more capable forecasters focuses on the foundation models (Bommasani et al., 2021), which address data-scarce scenarios by pre-training. Recent models aim to generalize on out-of-distribution data, achieving training-free overheads like statistical methods while enjoying the capacity as deep models. Another aspect is the shift from point to probabilistic forecasting (Woo et al., 2024; Ansari et al., 2024), which greatly improves the experience across many use cases by addressing forecasting uncertainty. While parametric densities can be easily incorporated into task-specific training, they can be overwhelmed by the heterogeneity of large-scale corpora. Applying this paradigm to pre-train foundation models is likely to result in mode collapse, manifested as over-smooth predictions given by the pre-trained model. In this work, we formally introduce generative time series foundation models, which naturally address the uncertainty in forecasting.
2.2 Time Series Foundation Models
Recent research has concentrated on building versatile large time series models (Liang et al., 2024). With the advance made in large language models, Transformer has become the dominant architecture. Several works adapt Transformers to address the unique 2D-dimensionality and heterogeneity of time series (Woo et al., 2024; Liu et al., 2024a). Specifically, our work delves into tokenization and optimization. Models such as TimesFM (Das et al., 2023b), Timer (Liu et al., 2024a, b), and Time-MoE (Shi et al., 2024b) embed continuous values and fit unimodal distributions via MSE or quantile loss (Wen et al., 2017). However, such prior loss may result in mode collapse in the of pre-training foundation models, and deterministic outcomes often fail to satisfy the requirement of decision-making. Based on continuous tokenization, Moirai (Woo et al., 2024) is a probabilistic model learning a mixture of distributions, but this prior can still struggle to fit complex distributions. Inspired by language modeling for scalable pre-training, Chronos (Ansari et al., 2024) discretizes series by scaling and quantization, learning more flexible categorical distribution by cross-entropy. Still, discrete tokenizer is limited to point level and sensitive to quantization interval, which is replaced by patch embedding and quantile head in subsequent work. Unlike before, we tame Transformer as native time series foundation model, learning flexible distributions without discrete tokenization.
2.3 Generative Modeling for Time Series
To address distributional heterogeneity during pre-training, generative modeling has become a focal point in the development of foundation models (Zhao et al., 2023; Liu et al., 2024c). While this direction for time series mostly focused on time series generation (Tashiro et al., 2021) and task-specific forecasters (Rasul et al., 2021; Shen & Kwok, 2023; Kollovieh et al., 2024), generative modeling for time series foundation models is hardly explored. With the comparable flexibility in distribution learning as language modeling, diffusion denoising (Sohl-Dickstein et al., 2015) and flow-matching (Lipman et al., 2022) have gained increasing prevalence in continuous-valued modalities (Lipman et al., 2024). Compared with diffusion denoising models, flow-matching provides a simple yet efficient framework. With fewer steps involved in the forward and reverse processes, large models based on flow-matching have shown superior performance in image generation (Esser et al., 2024).
Despite the connection in value continuity, generating images and future time series are fundamentally different tasks due to the autoregressive property of forecasting. Our proposed TimeFlow Loss is designed for autoregressive models to conduct conditional generation, which is a parameterized loss function (Zhang et al., 2018) for arbitrary distributions that enhances representation learning of foundation models.
3 Preliminaries
3.1 Flow-Matching
The goal of generative modeling is to learn the underlying probability distribution that generates the data. The framework of flow-matching transforms a sample $\mathbf{x}_{0}\sim p_{0}$ drawn from a source distribution into a sample $\mathbf{x}_{1}\sim p_{1}$ drawn from a target distribution. The transformation is continuous in time. For $d$-dimensional distributions, it is defined by a time-dependent velocity field $u_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ , which is the solution of the ordinary differential equation (ODE):
$\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}(\mathbf{x})=u_{t}\big{(}\psi_{t}(\mathbf{x})\big{)}\ \text{and}\ \psi_{0}(\mathbf{x})=\mathbf{x}.$
The velocity field $u_{t}$ determines a flow $\psi_{t}$. For all $t\in[0,1]$, $\psi_{t}$ generates the probability path $p_{t}$ that interpolates $p_{0}$ and $p_{1}$, i.e., $\mathbf{x}_{t}=\psi_{t}\left(\mathbf{x}_{0}\right)\sim p_{t}$ for $\mathbf{x}_{0}\sim p_{0}$. The implementation of flow-matching is to train a network $u^{\theta}_{t}$ parametrized by $\theta$ to fit the velocity field $u_{t}$, which is a regression-based task formulated as the Flow-Matching objective:
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{t}}\left\|u^{\theta}_{t}(\mathbf{x}_{t})-u_{t}(\mathbf{x}_{t})\right\|^{2}.$
Furthermore, Lipman et al. (2022) proved the equivalence of optimizing the Conditional Flow-Matching objective:
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{t},\mathbf{x}_{1}}\left\|u^{\theta}_{t}(\mathbf{x}_{t})-u_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})\right\|^{2}.$
Leveraging the conditional optimal-transport (linear) path and a source Gaussian, the objective can be formulated as:
$\mathcal{L}^{\mathrm{Gauss}}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,\epsilon,\mathbf{x}_{1}}\left\|u^{\theta}_{t}(\mathbf{x}_{t})-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|^{2}.$ (1)
where $t\sim\mathcal{U}[0,1],\mathbf{x}_{0}\sim\mathcal{N}(0,1)$ and $\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\epsilon$.
Consequently, we can train a generative network on given samples from the target distribution, and generate new samples by applying a push-forward process on samples drawn from a simple source Gaussian distribution:
$\mathbf{x}_{t+\Delta t}-\mathbf{x}_{t}=u^{\theta}_{t}(\mathbf{x}_{t})\Delta t,\ \mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\ t\in[0,1].$ (2)
Figure 2: Overall architecture of Sundial. The input time series is divided into patch tokens, which are embedded from original continuous values. The patch embeddings are fed into a decoder-only Transformer, a stable and speedup version that learns token representations via causal self-attention. The model is optimized using our TimeFlow Loss, a parameterized loss function that models per-token probability distribution conditioned on the learned representations, and generates multiple plausible predictions under the flow-matching framework.
3.2 Generative Models for Probabilistic Forecasting
Given a historical observation $x_{1:t}=\{x_{1},...,x_{t}\}$, the target of time series forecasting is to predict future time series $x_{t+1:t+f}=\{x_{t+1},\dots,x_{t+f}\}$. The task can be generally formulated as $p(x_{t+1:t+f}|\mathbf{h}_{t})$, where $\mathbf{h}_{t}=f_{\phi}(x_{1:t})$ is the learned representation from a deep model $f_{\phi}$. In probabilistic forecasting, explicit optimization objectives are utilized to predict the statistics of future series, e.g., MSE or quantiles, which have specified $p$ as a prior distribution. While using one parametric density generally fits well on a small amount of data, it can be the major bottleneck for scaling time series foundation models. Inspired by the success of large generative models (Rombach et al., 2022; OpenAI, 2023; Esser et al., 2024), we introduce generative modeling to realize probabilistic forecasting:
$p(x_{t+1:t+f}|\mathbf{h}_{t})=g_{\theta}\big{(}f_{\phi}({x_{1:t}})\big{)}.$ (3)
$g_{\theta}$ is a small trainable generative network conditioned on the learned representations of $f_{\phi}$. Both of them are jointly optimized. While the generative model automatically learns the distribution, it supports probabilistic forecasting by generating raw predictions and calculating their statistics. The idea is conceptually related to conformal prediction (Angelopoulos & Bates, 2021) but enjoys flexibility in outputs.
4 Approach
In this work, we conduct a univariate pre-training paradigm, which adopts Channel-Independence (Nie et al., 2022) on multivariate data. To mitigate value range discrepancy in time series, we conduct normalization on time series individually per variable. Afterwards, we sample varying-length training samples with the maximum context length of $2880$. As a foundation model, Sundial is required to predict on out-of-distribution series with varied lengths during inference.
4.1 Sundial
As shown in Figure 2, our Sundial models consist of three parts: (1) time series tokenization, including a context-level re-normalization and a patch embedding that addresses any-length time series, (2) a Transformer backbone that learns the per-token representation of time series, and (3) TimeFlow Loss, a parameterized loss function to model the per-token distribution and generate raw series during inference. Sundial is designed to operate on continuous-valued time series and facilitates scalable pre-training on heterogeneous distribution. Additionally, Sundial presents a new approach to probabilistic forecasting via generative modeling.
4.1.1 Time Series Tokenization
Re-Normalization
We adopt re-normalization (Liu et al., 2022), a non-parametric two-stage instance normalization conducted within each sample. While it is initially proposed to mitigate non-stationarity of time series, it addresses the value discrepancies and temporal distribution shifts, improving generalizability for zero-shot forecasting.
Patch Embedding
Given a univariate time series $\mathbf{X}=\{x_{1},\dots,x_{T}\}$, it is divided into patches $\mathbf{x}_{i}=x_{1+(i-1)P:iP}$ with the length of $P$. To address non-divisible length, we pad the input at the beginning and use a binary mask $\mathbf{m}_{i}\in\mathbb{R}^{P}$ for each patch to indicate the padded position. It will lead to $N=\lceil T/P\rceil$ such input tokens. Subsequently, we use a shared MLP $:\mathbb{R}^{2P}\mapsto\mathbb{R}^{D}$ to embed all patch tokens:
$\mathbf{h}_{i}=\operatorname{PatchEmbed}\big{(}\operatorname{Concat}(\mathbf{x}_{i},\mathbf{m}_{i})\big{)},$ (4)
where $\mathbf{h}_{i}\in\mathbb{R}^{D}$ and $D$ is the dimension of token embedding. Unlike point-level quantization (Ansari et al., 2024), we reserve original values without discrete quantization while reducing the number of tokens feeding to the Transformer.
4.1.2 Transformer Backbone
Given $N$ token embeddings $\{\mathbf{h}_{i}\}$, we adopt three crucial adaptations on a decoder-only Transformer to obtain per-token representations aggregated from all previous tokens. First, we adapt Pre-LN (Xiong et al., 2020) to improve pre-training stability. Second, we leverage a causal self-attention mechanism with RoPE (Su et al., 2024) that introduces the position information of patch tokens. It can be formulated as follows (the layer index is omitted for simplicity):
$\mathcal{A}_{ij}$ $=\mathbf{h}_{i}^{\top}\mathbf{W}_{\mathbf{q}}\mathbf{R}_{\Theta,i-j}\mathbf{W}_{\mathbf{k}}^{\top}\mathbf{h}_{j},$ (5)
$\operatorname{Attention}(\mathbf{H})$ $=\operatorname{Softmax}\left(\frac{\operatorname{Mask}(\mathcal{A})}{\sqrt{d}}\right)\mathbf{H}\mathbf{W}_{\mathbf{v}},$
where $\mathbf{W}_{\mathbf{q}},\mathbf{W}_{\mathbf{k}},\mathbf{W}_{\mathbf{v}}\in\mathbb{R}^{D\times d}$ project token embeddings $\mathbf{H}=\{\mathbf{h}_{i}\}$ into $d$-dimensional queries, keys, and values. $\mathbf{R}_{\Theta,t}\in\mathbb{R}^{d\times d}$ is the rotary matrix with rotation degree $(t\cdot\Theta)$. Lastly, we implement FlashAttention (Dao et al., 2022) and KV Cache (Pope et al., 2023), since these enhancements for deployment are increasingly emphasized in large foundation models (Shoeybi et al., 2019; Rasley et al., 2020).
4.1.3 TimeFlow Loss
Given representations $\{\mathbf{h}_{i}\}$ extracted by the last layer of the Transformer, we aim to generate length-$F$ predictions $\widehat{\mathbf{y}}_{i}=\widehat{x}_{1+iP,F+iP}$ at each position $i$ via our autoregressive model. While a small input patch size can accommodate more high-frequency data, a large output patch size can empirically lead to better results (Das et al., 2023b). Therefore, we adopt $F>P$ for multi-patch predictions in our models.
Based on Equations 1 and 3, we formulate a new generative forecasting conditioned on a sequential representation $\mathbf{h}_{i}$:
$\mathcal{L}(\theta,\mathbf{h}_{i})=\mathbb{E}_{t,\epsilon,\mathbf{y}_{i}}\left\|u^{\theta}_{t}\big{(}\mathbf{y}_{i}^{(t)}|\mathbf{h}_{i}\big{)}-\big{(}\mathbf{y}_{i}-\mathbf{y}_{i}^{(0)}\big{)}\right\|^{2}.$ (6)
where $\mathbf{y}_{i}\in\mathbb{R}^{F}$ is the groundtruth value and $\mathbf{y}_{i}^{(0)}$ is a $d$-dimensional Gaussian noise, $t$ is sampled from $\mathcal{U}[0,1]$, and $\mathbf{y}_{i}^{(t)}=t\mathbf{y}_{i}+(1-t)\mathbf{y}_{i}^{(0)}$ constructed by the conditional optimal-transport path. It is important to note that the conditional representation $\mathbf{h}_{i}$ differs from the conditional path and the conditional source distribution. Instead, $\mathbf{h}_{i}$ is a condition of position $i$, also a time-invariant condition of the whole flow-matching process $t\in[0,1]$. Technically, we implement the flow-matching network by a small MLP:
$u^{\theta}_{t}\big{(}\mathbf{y}_{i}^{(t)}|\mathbf{h}_{i}\big{)}=\operatorname{FM-Net}\big{(}\mathbf{y}_{i}^{(t)},t,\mathbf{h}_{i}\big{)}.$ (7)
The training process involves sampling the noised $\mathbf{y}_{i}^{(t)}$, and jointly input it with $t$. The condition $\mathbf{h}_{i}$ is integrated into the flow-matching network via AdaLN (Peebles & Xie, 2023). TimeFlow Loss for autoregressive models is formulated as:
$\mathcal{L}_{\mathrm{TimeFlow}}=\sum_{i=1}^{N}\left\|\operatorname{FM-Net}\big{(}\mathbf{y}_{i}^{(t)},t,\mathbf{h}_{i}\big{)}-\big{(}\mathbf{y}_{i}-\mathbf{y}_{i}^{(0)}\big{)}\right\|^{2}.$ (8)
Inference
Based on Equation 2, the push-forward process conditioned on a learned representation $\mathbf{h}_{i}$ is formulated as
$\mathbf{y}_{i}^{(t+\Delta t)}=\mathbf{y}_{i}^{(t)}+u^{\theta}_{t}\big{(}\mathbf{y}_{i}^{(t)}|\mathbf{h}_{i}\big{)}\Delta t.$ (9)
Technically, we adopt a $K$-step uniform trajectory, and set $\Delta t=1/K$. The sampling is done via starting from an initial Gaussian noise and advancing with the velocity generated by the trained $\operatorname{FM-Net}$ iteratively, as shown in Algorithm 1.
Algorithm 1 TimeFlow Loss: Sampling
0:  condition $\mathbf{h}_{i}\in\mathbb{R}^{D}$, path steps $K$.
1:  Sample initial noise $\widehat{\mathbf{y}}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
2:  $\Delta t=1/K$
3:  $\textbf{for}\ k\ \textbf{in}\ \{0,1\dots,K-1\}\ \textbf{do}$
4:  for $\widehat{\mathbf{y}}_{i}\leftarrow\widehat{\mathbf{y}}_{i}+\operatorname{FM-Net}\big{(}\widehat{\mathbf{y}}_{i},k\Delta t,\mathbf{h}_{i}\big{)}\Delta t$
5:  end for
6:  Return: $\widehat{\mathbf{y}}_{i}$
This procedure generates a predicted sample $\widehat{\mathbf{y}}_{i}$ at position $i$. To facilitate probabilistic forecasting, the procedure can be repeated using various initial noises, thereby enabling the computation of various statistics such as the median and quantiles from a set of generated probable predictions.
4.2 TimeBench
We collected and curated TimeBench, which comprises over 1 trillion time points from various sources, as shown in Figure 3. Several datasets originate from research teams (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024a, b). While most datasets are collected from real-world records, a small portion (0.05%) is generated synthetically to enhance pattern diversity, following KernelSynth proposed by Ansari et al. (2024). We also leverage substantial meteorological data (Hersbach et al., 2020) because of the predictability of weather systems. Data of different frequencies implies common and comprehensive temporal dynamics.
Figure 3: Ratios of data sources in TimeBench, the pre-training corpora of Sundial. Detailed statistics are provide in Table 4.
5 Experiments
We extensively evaluate Sundial on zero-shot forecasting benchmarks (Section 5.1) and investigate the scaling behavior across different model sizes (Section 5.2). We validate the effectiveness of TimeFlow compared to other training objectives (Section 5.3). We discuss the performance/speed trade-off during inference (Section 5.4) and conduct model adaptation to justify transferability of Sundial (Section 5.5).
Figure 4: Evaluation summary of Sundial.
Table 1: Zero-shot forecasting results of time series foundation models on long-term forecasting datasets (Time-Series-Library) (Wu et al., 2022). Corresponding prediction lengths include $\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. Averaged results of four prediction lengths are reported here. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets. Results of baseline models are officially reported by Shi et al. (2024b). Datasets in pre-training are not evaluated on corresponding models, which are denoted by the dash (--). Full results under all prediction lengths are provided in Table 7.
Models
$\textbf{Sundial}_{\textit{Small}}$
$\textbf{Sundial}_{\textit{Base}}$
$\textbf{Sundial}_{\textit{Large}}$
[…]
Timer
$\textbf{Moirai}_{\textit{Base}}$
$\textbf{Moirai}_{\textit{Large}}$
$\textbf{Chronos}_{\textit{Base}}$
$\textbf{Chronos}_{\textit{Large}}$
TimesFM
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.354
0.388
0.336 0.377
0.331
0.369
0.394
0.415
0.376
0.405
0.356
0.391
0.373
0.392
0.406
0.385
0.422
0.391
0.645
0.500
0.555
0.465
0.433
0.418
ETTm2
0.265
0.324
0.258 0.320
0.254
0.315
0.317
0.365
0.316
0.361
0.288
0.344
0.273
0.336
0.311
0.337
0.329
0.343
0.310
0.350
0.295
0.338
0.328
0.346
ETTh1
0.390
0.418
0.411
0.434
0.395
0.420
0.400
0.424
0.394
0.419
0.412
0.426
0.404
0.417
0.417
0.419
0.480
0.439
0.591
0.468
0.588
0.466
0.473
0.443
ETTh2
0.340
0.387
0.333
0.387
0.334
0.387
0.366
0.404
0.405
0.415
0.371
0.399
0.347
0.388
0.362
0.382
0.367
0.377
0.405
0.410
0.455
0.427
0.392
0.406
ECL
0.169 0.265 0.169 0.265
0.166
0.262
0.174
0.278
0.187
0.274
0.186
0.270
0.214
0.278
0.204
0.273
Weather
0.233
0.271 0.234
0.270
0.238
0.275
0.265
0.297
0.270
0.300
0.256
0.288
0.256
0.294
0.287
0.281
0.264
0.273
0.292
0.315
0.279
0.306
$1^{\text{st}}$ Count
7
2
8
5
16
16
0
1
0
0
2
1
1
3
0
2
0
6
0
0
0
0
0
0
Table 2: Aggregated performance on GIFT-Eval, which comprises $23$ datasets characterized by a variety of frequencies, number of variates, and prediction lengths. We evaluate zero-shot forecasting performance using Sundial (Base). A lower MASE or CRPS indicates a better prediction. Rank assigns a numerical ranking of all $97$ configurations. Results of baselines are officially reported by Aksu et al. (2024).
Type Statistical Methods Task-Specific Models (Superwised) Time Series Foundation Models (Zero-Shot)
Model Naïve Seasonal Auto Auto DeepAR TiDE NBEATS PTST. iTrans. TimesFM VisionTS Chronos Moirai Sundial
Naïve ARIMA Theta (2020) (2023a) (2019) (2022) (2023) (2023b) (2024) (2024) (2024) (Ours)
MASE 1.260 1.000 0.964 0.978 1.206 0.980 0.842 0.762 0.802 0.967 0.775 0.786 0.809 0.727
CRPS 1.383 1.000 0.770 1.051 0.721 0.652 0.689 0.496 0.524 0.575 0.638 0.551 0.515 0.505
Rank 23.392 21.598 17.464 19.928 15.381 14.856 17.309 7.680 8.580 11.443 16.474 11.247 7.845 9.536
5.1 Time Series Forecasting
In this section, we focus on zero-shot forecasting, we compare Sundial with advanced time series foundation models on various benchmarks, including (1) point forecasting: we adopt the long-term forecasting benchmark (Wu et al., 2022), which assesses the performance under different forecasting horizons using MSE and MAE; (2) probabilistic forecasting: we experiment on GIFT-Eval (Aksu et al., 2024) and FEV leaderboard (Ansari et al., 2024), following their official evaluation suite and assessing point (MASE) and probabilistic (CRPS and WQL) metrics. All evaluated datasets are excluded from the pre-training dataset. The model configurations are detailed in Table 5.
5.1.1 Point Forecasting
As shown in Table 1, Sundial consistently outperforms other advanced time series foundation models. Compared with the previous state-of-the-art model Time-MoE (Shi et al., 2024b), the Sundial family using fewer parameters achieves the average MSE reduction of $7.57\%$ and averaged MAE reduction of $4.71\%$. Notably, continuous tokenization allows our model to conduct patch-level forecasting with fewer autoregression steps, while Chronos using point-wise discrete tokenization may not be effective in long-term forecasting.
Figure 5: Model evaluation on the FEV leaderboard, which includes $27$ datasets not seen by Sundial. Baseline models can be categorized into statistical methods fitting on each time series, task-specific deep models trained on each dataset, and pre-trained foundation models. Pre-trained Models that have seen several datasets during pre-training are denoted as Pre-trained Models (Other). A lower MASE/WQL indicates a better result. Sundial makes probabilistic predictions using $20$ generated series, being consistent with Ansari et al. (2024).
5.1.2 Probabilistic Forecasting
Beyond point forecasting, Sundial possesses a unique generative capability for making probabilistic predictions. Following Ansari et al. (2024), we calculate the median and quantiles using $20$ generated raw predictions of Sundial. While several baseline models have been pre-trained by the consistent objective function for probabilistic evaluation, e.g., quantile loss for WQL, Sundial calculates these statistics for evaluation without any prior knowledge.
GIFT-Eval
Aggregated results are presented in Table 2. The benchmark evaluates performance from $23$ datasets and $13$ baseline models, encompassing statistical methods, task-specific models, and time series foundation models. Among supervised models and advanced foundation models, Sundial attains the first place in MASE and second place in CRPS on all unseen datasets. While the top PatchTST (Nie et al., 2022) is exhaustively trained and tweaked on each dataset, the zero-shot performance of Sundial highlights its simplicity and robustness on this comprehensive benchmark.
FEV Leaderboard
We evaluate our Sundial on the open leaderboard established by AutoGluon (Ansari et al., 2024), which includes $27$ datasets for probabilistic forecasting. As shown in Figure 5, the zero-shot forecasting performance of Sundial exceeds $60\%$ statistical methods and deep models that are superwisedly trained in distribution. While Sundial is ranked as the second zero-shot pre-trained models after Chronos, Sundial realizes 11.34×11.34$\times$11.34 × inference time speedup as shown in Figure 6. Based on native patching and multi-patch prediction, our inference time is near to N-BEATS.
Qualitative showcases for point and probabilistic forecasting are presented in Appendix D. Our model generates highly eventful and coherent temporal patterns with input series.
Figure 6: Inference time evaluation following Ansari et al. (2024), which is averaged from the FEV leaderboard. Computing resources of different models are marked. We plot the logarithmic x-axis.
5.2 Scalability
From Table 1, the larger Sundial model generally achieves better performance and more wins with the scaling of parameters. Beyond downstream performance, we delve into the utilization of model capacity. We plot training curves in Figure 7. Compared with Sundial (Small), the large version leads to totally $15.38\%$ reduction in the converged training objective, showing significant performance promotion on in-distribution time series forecasting.
Figure 7: Training curves on TimeBench of different model sizes.
5.3 TimeFlow Loss
Based on the flow-matching framework, TimeFlow Loss allows autoregressive models to learn and generate flexible distributions based on learned representations. To validate the effectiveness of this design, we implement two alternatives: (1) an MLP network and MSE Loss and (2) a parameterized training objective based on the denoising diffusion procedure (Li et al., 2024). We adopt the same parameterized network and Transformer backbone and pre-train them on TimeBench. Since the converged training loss is not comparable across different objective functions, we compare zero-shot performance in Table 3. Despite allowing for prediction generation, performance using diffusion-based objective is notably inferior to TimeFlow Loss.
Table 3: Zero-shot performance using different training objectives. We use the same model configuration and pre-training scale. Averaged MSE of four prediction lengths are reported here.
Objective
ETTm1
ETTm2
ETTh1
ETTh2
ECL
Weather
Avg.
TimeFlow
0.336
0.258
0.411
0.333
0.169
0.234
0.290
Diffusion
0.362
0.265
0.444
0.360
0.202
0.252
0.314
MSE
0.360
0.264
0.404
0.341
0.175
0.231
0.296
In addition to zero-shot performance, we provide showcases for quality evaluations in Appendix D.2. Pre-trained model optimized by the specific MSE Loss can only output a single prediction and the prediction is sometimes over-smooth due to mode collapse in large-scale pre-training. During pre-training, generative modeling can accommodate significantly different future variations even if their lookback series are similar. It benefits downstream tasks by generating multiple plausible predictions, which indicates various possibilities in time series forecasting, thereby facilitating the applicability for decision-making.
5.4 Model Inference
Generative modeling introduces flexibility to adjust predictions of pre-trained models during inference. Even with the naïve strategy to generate predictions, i.e., sampling different noise from a standard Gaussian distribution, there are two configurations to control the quality of predictions: (1) the number of sampled predictions to calculate statistics and (2) sampling steps specified in flow-matching. We present results with varied configurations in Figure 8.
Figure 8: We show the MASE (left) and WQL (right) on FEV w.r.t. the number of generated raw predictions (top) and the steps to sample a prediction (down). More predictions or more sampling steps generally achieve better probabilistic metrics.
The top two figures conform to the central limit theorem. Generating more samples consistently leads to more precise estimation of statistics and better results. The bottom two figures indicate that using fine-grained steps during the push-forward process generally leads to good predictions.
These observations reveal the trade-off between inference speed and performance, which provides flexibility for various use cases requiring different certainty in predictions. Note that using different configurations does not require retraining the model. In our experiments, we consistently adopt the choice of sampling $20$ predictions with each generated by $50$ steps, achieving comparable inference time as task-specific deep models as shown in Figure 6. Advanced strategies of sampling and post-processing of raw prediction leave interesting directions for future exploration.
5.5 Model Adaptation
Figure 9: Performance on the FEV leaderboard, including (1) training Sundial from scratch on all datasets from the FEV leaderboard, (2) zero-shot forecasting using pre-trained Sundial, and (3) fine-tuning once on all datasets from the FEV leaderboard.
Inspired by the prevalence of instruction tuning (Wei et al., 2021) that adapts foundation models on a collection of tasks. We fine-tune pre-trained Sundial (Base) on the FEV leaderboard, including short-term tasks with different prediction lengths. Our model is tuned once on all aggregated datasets. We evaluate the performance on unseen test splits (Figure 9). We observe that the performance can be further improved compared to zero-shot forecasting. Furthermore, training from scratch on aggregated datasets results in inferior performance, implying knowledge transfer in pre-trained models.
6 Conclusion
In this work, we collect and curate TimeBench, a trillion-scale time series dataset for building time series foundation models, which can benefit the research community. Towards time series foundation models, we delve into tokenization and optimization, presenting contributions in two aspects. First, we demonstrate that continuous tokenization, such as patch series, can be more effective and efficient for the time series modality, and generative modeling presents a native approach for learning on continuous-valued time series. Second, we propose a novel training objective to accommodate heterogeneous time series distribution while endowing autoregressive models with an inherent capability to sample from non-categorical distribution. Our pre-trained Sundial models make substantial advances on best-recognized forecasting leaderboards. We hope this work can inspire future paradigms for pre-training time series foundation models and enhance their applicability to real-world scenarios.
Appendix A Dataset Statistics
Large-scale datasets are of paramount importance for pre-training foundation models. Recent research has contributed significant time series datasets (Das et al., 2023b; Liu et al., 2024b; Shi et al., 2024b). While the scaling law of time series foundation models has been explored in the recent work (Shi et al., 2024a), the pre-training scale remains relatively limited. Given the heterogeneity of time series compared to other modalities, it raises the question of whether it is feasible to learn from enormous series. To address the question, we curated TimeBench including 1 trillion time points from various domains.
The statistical details of TimeBench are summarized in Table 4. In addition to open-source datasets from research teams on time series foundation models (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024b, a), we collected substantial real-world time series from various domains such as finance, IoT, meteorology (Muñoz-Sabater et al., 2021), and healthcare (Goldberger et al., 2000). These resources enable us to construct large-scale time-series corpora exceeding 1 trillion time points. The corpora include highly credible and predictable data with a wide range of frequencies, lengths, and number of variates, providing comprehensive temporal dynamics and variation patterns to facilitate downstream applications. To prevent data leakage, we exclude all datasets evaluated in Section 5.1 to make sure that Sundial conducts zero-shot forecasting.
Table 4: Key statistics of TimeBench, the pre-training dataset of Sundial, which encompasses various sources.
Source Chronos ECG Finance IoT LOSTA Synthetic ERA5 3h ERA 12h ERA5 Daily
ERA5 Weekly
ERA5 Monthly
ERA5 Quarterly
Total
(2024) (2000) (Ours) (Ours) (2024) (2024) (2021) (2021) (2021) (2021) (2021) (2021)
# Pts. 94B 48B 10.5B 5.8B 230B 0.5B 129B 32B 406B 58B 13.5B 4.5B 1032B
% 9.11 % 4.65 % 1.02 % 0.56 % 22.29 % 0.05 % 12.50 % 3.10 % 39.35 % 5.62 % 1.31 % 0.44 % 100%
Appendix B Implementation Details
All experiments are implemented using PyTorch (Paszke et al., 2019) and executed on NVIDIA A100 GPUs. We employ the AdamW optimizer (Kingma & Ba, 2014) for model optimization. We adopt channel independence (Nie et al., 2022) for univariate pre-training. During training, data from different domains is sampled according to a predefined ratio to balance the data volume across domains and ensure diversity in the training data. We implement a global shuffle strategy by loading time series into a standard parquet format. We use variable-wise normalization to unify the scope of values.
On the FEV leaderboard (Ansari et al., 2024), which consists of short-term forecasting datasets with a maximum prediction length of $56$, we train Sundial models by TimeFlow Loss with the prediction length of $F=16$. For the point forecasting (Wu et al., 2022) and GIFT-Eval (Aksu et al., 2024), which consist of forecasting datasets with a prediction length ranging from $6$ to $900$, we train Sundial models by TimeFlow Loss with the prediction length of $F=720$. For the required prediction length less than the model prediction length, we truncate the output generated by Sundial. For the required length more than the prediction horizon, we conduct rolling forecasting. Following the generative forecaster Chronos (Ansari et al., 2024), we sample $20$ raw predictions with each generated by $50$ sampling steps to calculate metrics for evaluation, including MASE, CRPS, and WQL. Detailed configurations of Sundial in different sizes are provided in Table 5. We provide a model summary in Table 6, which includes more counterparts and summarizes several essential aspects of time series foundation models.
Table 5: Model configurations of the Sundial family.
Model
Patch Size
Context Length
Prediction Length
Layers
Dimension
MHA Heads
TimeFlow
Total Parameters
[…]
$\textbf{Sundial}_{\textit{Small}}$ $16$ $2880$$\{16,720\}$ 6 $(512,2048)$ $8$ $(512,3)$ $32$M
$\textbf{Sundial}_{\textit{Base}}$ $16$ $2880$$\{16,720\}$ 12 $(768,3072)$ $12$ $(768,3)$ $128$M
$\textbf{Sundial}_{\textit{Large}}$ $16$ $2880$$\{16,720\}$ 24 $(1024,4096)$ $16$ $(1024,6)$ $444$M
  • $\ast$ $D$ is the embedding dimension of Transformer. $D_{\text{ff}}$ is the hidden dimension of FFN. $D_{\text{tf}}$ is the hidden dimension of the flow-matching network. $L$ is the layer number of Transformer. $L_{\text{tf}}$ is the layer number of the flow-matching network.
Table 6: Comparison of time series foundation models. Architecture denotes the Transformer category. Model Size presents parameter counts of different model sizes. Pre-training Scale measures pre-training datasets in time points. Token Level presents the graininess of time series tokens. Tokenization denotes what kind of values are embedded from time series. Context Length means the input length supported by the model. Probabilistic means generating multiple probable predictions, which is the opposite of deterministic forecasters.
Method Sundial Time-MoE Timer Moirai MOMENT LLMTime Chronos Lag-Llama TimesFM
(Ours) (2024b) (2024a) (2024) (2024) (2024) (2024) (2023) (2023b)
Architecture Decoder Decoder Decoder Encoder Encoder Decoder EncDec Decoder Decoder
Model Size 32M 113M 29M 14M 40M - 46M 200M 17M
128M 453M 50M 91M 125M 200M 70M
 [^4] M 2.4B 67M 311M 385M 710M 200M
Pre-training Scale 1034B 300B 231B 231B 1.13B - 84B 0.36B 100B
Token Level Patch Point Patch Patch Patch Point Point Point Patch
Tokenization
Continuous
Continuous
Continuous
Continuous
Continuous
Discrete
Discrete
Continuous
Continuous
Context Length $\leq$2880 $\leq$4096 $\leq$1440 $\leq$5000 = 512 - $\leq$512 $\leq$1024 $\leq$512
Probabilistic True False False True False True True True False
Appendix C Supplementary Results
C.1 Zero-Shot Results of Point Forecasting
Table 7 provides full zero-shot results in Time-Series-Library forecasting benchmark (Wu et al., 2022), including prediction horizons in $\{96,192,336,720\}$. We build Sundial with different model sizes with configurations in Table 5. The context length is set as $2880$. We truncate the model’s predictions for tasks requiring a prediction length less than $F=720$.
We compare most advanced time series foundation models based on their official results, including Time-MoE (Shi et al., 2024b), Timer (Liu et al., 2024a, b), Moirai (Woo et al., 2024), TimesFM (Das et al., 2023b), and Chronos (Ansari et al., 2024). We conduct zero-shot evaluations on datasets that are not included during the pre-training of corresponding models. For each of the evaluated model, we use their maximum input length during inference. The metric (MSE/MAE) is calculated from all predicted windows in the test split of each dataset.
C.2 Zero-Shot Results on GIFT-Eval and FEV Leaderboard
We evaluate our models on GIFT-Eval, a benchmark designed to comprehensively assess forecasting performance across diverse time series. GIFT-Eval includes $23$ datasets covering $144,000$ time series and $177$ million data points, which constitute a total of $97$ forecasting configurations. We use the official evaluation suite established by the research team of SalesForce and report aggregated results in Table 2. We evaluate the probabilistic forecasting performance on the FEV leaderboard, which was originally proposed by Ansari et al. (2024) and established by AutoGluon, which comprises $27$ datasets for zero-shot evaluation. We report aggregated metrics in Figure 5 and assess the inference time in Figure 6. We will release the detailed results by submitting Sundial to their open benchmark in the future.
Appendix D Showcases
D.1 Showcases of Sundial
We present zero-shot forecasting showcases on all the datasets from FEV (Ansari et al., 2024) in Figure 10-11, and long-term forecasting datasets (Wu et al., 2022) in Figure 12. By generating multiple predictions with different initial noise, we use the raw prediction to calculate the median and plot the $80\%$ prediction interval.
Table 7: Zero-shot forecasting results of time series foundation models on long-term forecasting datasets (Wu et al., 2022). A lower MSE or MAE indicates a better prediction. Averaged results of four prediction lengths are reported here. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets. Results of baseline models are officially reported by Shi et al. (2024b). Datasets for pre-training are not evaluated on corresponding models, which are denoted by the dash (--).
Models
$\textbf{Sundial}_{\textit{Small}}$
$\textbf{Sundial}_{\textit{Base}}$
$\textbf{Sundial}_{\textit{Large}}$
[…]
Timer
$\textbf{Moirai}_{\textit{Base}}$
$\textbf{Moirai}_{\textit{Large}}$
$\textbf{Chronos}_{\textit{Base}}$
$\textbf{Chronos}_{\textit{Large}}$
TimesFM
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.292
0.342
0.280 0.334
0.273
0.329
0.338
0.368
0.309
0.357
0.281
0.341
0.317
0.356
0.363
0.356
0.380
0.361
0.454
0.408
0.457
0.403
0.361
0.370
192
0.337
0.376
0.321
0.366
0.312
0.357
0.353
0.388
0.346
0.381
0.305
0.358
0.358
0.381
0.388
0.375
0.412
0.383
0.567
0.477
0.530
0.450
0.414
0.405
336
0.370
0.401
0.350 0.389
0.343
0.378
0.381
0.413
0.373
0.408
0.369
0.395
0.386
0.401
0.416
0.392
0.436
0.400
0.662
0.525
0.577
0.481
0.445
0.429
720
0.418
0.433
0.394
0.418 0.397
0.413
0.504
0.493
0.475
0.477
0.469
0.472
0.430
0.431
0.460
0.418
0.462
0.420
0.900
0.591
0.660
0.526
0.512
0.471
Avg
0.354
0.388
0.336 0.377
0.331
0.369
0.394
0.415
0.376
0.405
0.356
0.391
0.373
0.392
0.406
0.385
0.422
0.391
0.645
0.500
0.555
0.465
0.433
0.418
ETTm2
96
0.178
0.260
0.170
0.256 0.172
0.255
0.201
0.291
0.197
0.286
0.198
0.288
0.189
0.277
0.205
0.273
0.211
0.274
0.199
0.274
0.197
0.271
0.202
0.270
192
0.235
0.304
0.229 0.300
0.227
0.296
0.258
0.334
0.250
0.322
0.235
0.312
0.241
0.315
0.275
0.316
0.281
0.318
0.261
0.322
0.254
0.314
0.289
0.321
336
0.287
0.342
0.281 0.337
0.275
0.331
0.324
0.373
0.337
0.375
0.293
0.348
0.286
0.348
0.329
0.350
0.341
0.355
0.326
0.366
0.313
0.353
0.360
0.366
720
0.360
0.390
0.351 0.387
0.343
0.378
0.488
0.464
0.480
0.461
0.427
0.428
0.375
0.402
0.437
0.411
0.485
0.428
0.455
0.439
0.416
0.415
0.462
0.430
Avg
0.265
0.324
0.258 0.320
0.254
0.315
0.317
0.365
0.316
0.361
0.288
0.344
0.273
0.336
0.311
0.337
0.329
0.343
0.310
0.350
0.295
0.338
0.328
0.346
ETTh1
96
0.341
0.381
0.348
0.385
0.346
0.383
0.357
0.381
0.350
0.382
0.349
0.379
0.369
0.391
0.376
0.392
0.381
0.388
0.440
0.393
0.441
0.390
0.414
0.404
192
0.381
0.408
0.393
0.418
0.386
0.410
0.384
0.404
0.388
0.412
0.395
0.413
0.405
0.413
0.412
0.413
0.434
0.415
0.492
0.426
0.502
0.524
0.465
0.434
336
0.405
0.424
0.422
0.440
0.410
0.426
0.411
0.434
0.411
0.430
0.447
0.453
0.418
0.423
0.433
0.428
0.485
0.445
0.550
0.462
0.576
0.467
0.503
0.456
720
0.433
0.458
0.481
0.493
0.438
0.459
0.449
0.477
0.427
0.455
0.457
0.462
0.423
0.441
0.447
0.444
0.611
0.510
0.882
0.591
0.835
0.583
0.511
0.481
Avg
0.390
0.418
0.411
0.434
0.395
0.420
0.400
0.424
0.394
0.419
0.412
0.426
0.404
0.417
0.417
0.419
0.480
0.439
0.591
0.468
0.588
0.466
0.473
0.443
ETTh2
96
0.272
0.332 0.271
0.333
0.269
0.330
0.305
0.359
0.302
0.354
0.292
0.352
0.283
0.342
0.294
0.330
0.296
0.330
0.308
0.343
0.320
0.345
0.315
0.349
192
0.329
0.374
0.327
0.376
0.325
0.373
0.351
0.386
0.364
0.385
0.347
0.379
0.340
0.379
0.365
0.375
0.361
0.371
0.384
0.392
0.406
0.399
0.388
0.395
336
0.357 0.399
0.354
0.402
0.354
0.400
0.391
0.418
0.417
0.425
0.406
0.419
0.366
0.400
0.376
0.390
0.390
0.390
0.429
0.430
0.492
0.453
0.422
0.427
720
0.401
0.442
0.381
0.435
0.389
0.443
0.419
0.454
0.537
0.496
0.439
0.447
0.397
0.431
0.416
0.433
0.423
0.418
0.501
0.477
0.603
0.511
0.443
0.454
Avg
0.340
0.387
0.333
0.387
0.334
0.387
0.366
0.404
0.405
0.415
0.371
0.399
0.347
0.388
0.362
0.382
0.367
0.377
0.405
0.410
0.455
0.427
0.392
0.406
ECL
96
0.134
0.231
0.132 0.229
0.130
0.227
0.141
0.237
0.160
0.250
0.153
0.241
0.154
0.231
0.152
0.229
192
0.154
0.251
0.152 0.250
0.150
0.247
0.159
0.254
0.175
0.263
0.169
0.255
0.179
0.254
0.172
0.250
336
0.174
0.271
0.173 0.271
0.170
0.268
0.177
0.272
0.187
0.277
0.187
0.273
0.214
0.284
0.203
0.276
720
0.215
0.307
0.218
0.311
0.214
0.307
0.219
0.308
0.228
0.309
0.237
0.313
0.311
0.346
0.289
0.337
Avg
0.169 0.265 0.169 0.265
0.166
0.262
0.174
0.278
0.187
0.274
0.186
0.270
0.214
0.278
0.204
0.273
Weather
96
0.158 0.206
0.157
0.205
0.157
0.208
0.160
0.214
0.159
0.213
0.157
0.211
0.171
0.225
0.220
0.217
0.199
0.211
0.203
0.238
0.194
0.235
192
0.205
0.253
0.205
0.251
0.207
0.256
0.210
0.260
0.215
0.266
0.208
0.256
0.221
0.271
0.271
0.259
0.246
0.251
0.256
0.290
0.249
0.285
336
0.254 0.290
0.253
0.289
0.259
0.295
0.274
0.309
0.291
0.322
0.255
0.290
0.274
0.311
0.286
0.297
0.274
0.291
0.314
0.336
0.302
0.327
720
0.315
0.336
0.320
0.336
0.327
0.342
0.418
0.405
0.415
0.400
0.405
0.397
0.356
0.370
0.373
0.354
0.337
0.340
0.397
0.396
0.372
0.378
Avg
0.233
0.271 0.234
0.270
0.238
0.275
0.265
0.297
0.270
0.300
0.256
0.288
0.256
0.294
0.287
0.281
0.264
0.273
0.292
0.315
0.279
0.306
$1^{\text{st}}$ Count
7
2
8
5
16
16
0
1
0
0
2
1
1
3
0
2
0
6
0
0
0
0
0
0
  • $\ast$ Traffic (PEMS, ) is not evaluated because it is included in the pre-training datasets of these time series foundation models.
Figure 10: Showcases of zero-shot predictions from Sundial (Base) on the FEV leaderboard (Ansari et al., 2024).
Figure 11: Showcases of zero-shot predictions from Sundial (Base) on the FEV leaderboard (Ansari et al., 2024).
Figure 12: Showcases of zero-shot predictions from Sundial (Base) on long-term forecasting datasets (Wu et al., 2022).
D.2 Showcases of Pre-trained Generative Forecasters and Deterministic Forecasters
As we introduce generative modeling in time series foundation models, we compare zero-shot forecasting showcases from two types of models in Figure 13-14, including (1) Sundial pre-trained by TimeFlow: as a generative forecaster, it can predict various future possibilities based on the lookback series. (2) Adopting the same backbone and pre-training on TimeBench, a Transformer pre-trained by MSE Loss: as a deterministic forecaster, the model can only output one prediction. During pre-training, the unimodal Gaussian prior specified by MSE can be infeasible to handle large-scale pre-training, manifested as sometimes over-smooth predictions in downstream forecasting tasks. Therefore, we hope this work can inspire future paradigms for pre-training time series foundation models and enhance their applicability to real-world scenarios.
Figure 13: Showcases of Sundial (Left) and a counterpart Transformer pre-trained by MSE Loss (Right).
Figure 14: Showcases of Sundial (Left) and a counterpart Transformer pre-trained by MSE Loss (Right).
Appendix E Limitations
Our models represent an initial effort to incorporate generative modeling into time series foundation models, which enables pre-training on heterogeneous time series without specifying any prior distribution. This approach mitigates mode collapse in representation learning and generates a diverse range of probable predictions compared to previous deterministic forecasters. Despite significant progress, the Sundial family still faces limitations, which tend to generate conservative predictions, such as underestimating trends given a rising slope in context. This hallucination is also observed in Chronos (Ansari et al., 2024). We have yet to determine whether the underlying cause arises from the pre-training distribution or the generative paradigm. This situation may also indicate new opportunities during inference. As we only adopt a naïve sampling strategy that begins with random Gaussian noises, it leaves much room for future improvement in sampling strategy and post-processing.
Another aspect of future development lies in model adaptation. Sundial is pre-trained in a univariate approach to address the discrepancy in variate numbers, which prevents it from explicitly utilizing variate correlations or covariate information. As an increasing number of studies address 2D dimensionality (Liu et al., 2024a; Woo et al., 2024), multivariate pre-training is likely to be conducted for domain-specific time series foundation models. Lastly, while autoregressive models provide flexibility in the input context length, multiple steps of autoregression for long output lengths may still lead to oversmooth predictions and unreliable results. In addition to the forecasting principle emphasizing the importance of complete information (long context), we consider that the output length should be determined based on downstream predictability, where instruction tuning of time series foundation models could serve as a promising solution.
Appendix F Societal Impacts
F.1 Real-World Applications
In this work, we present a family of time series foundation models designed to facilitate zero-shot forecasting. Our models employ native tokenization for continuous-valued time series and incorporate a flexible training objective, proposed as TimeFlow Loss, to enable probabilistic forecasting. With an unprecedented distribution learning capability and a trillion-level pre-training scale, our models can be used directly or adapted for various forecasting scenarios, such as energy planning, device maintenance, and financial risk prevention. With multiple predictions that are highly coherent with the input series, our model enhances the reliability of decision-making and streamlines the forecasting pipeline for practitioners. This paper primarily focuses on scientific research and does not present any evident negative social impact.
F.2 Academic Research
We curate TimeBench, a trillion-level time series corpora for pre-training foundation models for time series analysis, which we believe will be beneficial to the research community. Technically, we propose a TimeFlow Loss to facilitate the learning of flexible next-patch distributions. Conditioned on the representations acquired by autoregressive Transformers, our model is endowed with a novel generative capability for probabilistic forecasting. It also enhances the representation learning of Transformers without the need for discrete tokenization. Through pre-training on an unprecedented dataset comprising one trillion time points, we identify subtle scalability bottlenecks that are not solely attributable to architectural design but are predominantly influenced by the training objectives of foundation models. The paradigm of generative modeling applied to autoregressive models may provide valuable insights for the development of continuous-valued foundation models.